3. A 2-kg steam water mixture at 1.0 MPaa is contained in an inflexible tank. Heat is

added until the pressure rises to 3.5 MPaa and the temperature to 400 o C. Determine:

a) Heat transferred (kJ), b) ΔH (kJ), c)ΔS (kJ/k)


1
Expert's answer
2021-05-18T21:46:01-0400

"m = 2kg\\
P_1 = 1.0\ MPa\\
P_2 = 3.5\ MPa\\
U_1 = 2401.41\ kJ/kg\\
U_2= 2926.4\ kJ/kg\\
T = 400°C"


Heat transferred = "m(U_2-U_1) ="

"2(2926.4-2401.41) = 2(524.99)"

"= 1049.98\ kJ"